The objective of a microscope has magnifying power of 20×, a numerical aperture of 0.4, a working distance (WD) of 4 mm (~f ). For a wavelength of light of 450 nm, what is the resolving power of this objective? What is the diameter of its lens?
Answer #1

The resolving power is given by

r=\frac{0.61\lambda }{na}

Given \lambda = 450\times 10^{-9}m , na=0.4

r=\frac{0.61\times 450\times 10^{-9}}{0.4}m

r=686.25nm

The diameter D will be given by

D=2\times f\times sin(na)

D=2\times WD\times sin(na)

D=2\times 4mm\times sin(0.4)=3.1153mm